# Practical hybrid decoding scheme for parity-encoded spin systems

### v2

@@ II.2 Parity encoding architecture and LDPC codes
We illustrate a PE architecture with concrete examples. The PE architecture corresponds to the classical ECC according to the following map: $\bar{z}_{ij}=\bar{Z}_{i}\oplus\bar{Z}_{j}$ for $1\leq i\leq j\leq k$. As a simple example, consider the case $K=4$ and assume that $\boldsymbol{\bar{Z}}=\left(\bar{Z}_{1},\bar{Z}_{2},\bar{Z}_{3},\bar{Z}_{4}\right)\in\left\{0,1\right\}^{K}$ and $\bar{\boldsymbol{z}}=\left(\bar{z}_{12},\bar{z}_{13},\bar{z}_{14},\bar{z}_{23},\bar{z}_{24},\bar{z}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$. Because each element of $\bar{\boldsymbol{z}}$ is the binary sum of two elements of $\boldsymbol{\bar{Z}}$, the generating matrix associated with this map is given by the $K\times\tbinom{K}{2}$ matrix:
@@
The connections between the variables $\bar{x}_{kl}$ and the weight-4 syndit romes $\bar{s}_{klmn}^{(4)}$ are depicted by a sparse bipartite graph shown in Fig.2(a). Similar connections between variables $\bar{x}_{kl}$ and the weight-3 syndromes $\bar{s}_{klm}^{(3)}$ are depicted in Fig.2(b). The number of elements in $\bar{\boldsymbol{x}}$ is $N_{v}=\tbinom{K}{2}$, while the number of elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ and $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ is $N_{c}=\tbinom{K-1}{2}$ and $N_{c}=\tbinom{K}{3}$, respectively. In the terminology of graph theory, the $N_{v}$ elements in $\bar{\boldsymbol{x}}$ constitute variable nodes (VNs) and the $N_{c}$ elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ or $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ constitute check nodes (CNs). The matrix $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$ has $N_{c}$ rows and $N_{v}$ columns, where each row $i$ represents a CN and each column $j$ represents a VN; If the entry $H_{ij}=1$, VN $j$ is connected to CN $i$ by an edge. Thus, each edge connecting VN and CN corresponds to an entry $1$ in the row and column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. Let $d_{c}$ be the number of $1$s in each row and $d_{v}$ be the number of $1$s in each column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. These numbers represent the number of edges connected to a VN and a CN, respectively, referred to as row and column weights. The matrix $\boldsymbol{H}^{\prime}$ is regular, where the row weight $d_{c}=3$ is common for all the CNs and the column weight $d_{v}=K-2$ is common for all the VNs. In contrast, the matrix $\boldsymbol{H}$ is irregular, and its weights depend on the associated nodes, which are at most 4.
@@ II.3 SLHZ model
In the remainder of this paper, we will discuss our arguments primarily in the language of spin glass. Following Sourlas, our argument relies on isomorphism between the additive Boolean group $\left(\left\{0,1\right\},\oplus\right)$ and the multiplicative Ising group $\left(\left\{\pm 1\right\},\cdot\right)$, where a binary variable $\bar{a}{}_{i}\in\left\{0,1\right\}$ maps to the spin variable $a_{i}=\left(-1\right)^{\bar{a}_{i}}\in\left\{\pm 1\right\}$ and the binary sum maps to the product by $a_{i}a_{j}=(-1)^{\bar{a}_{i}\oplus\bar{a}_{j}}\in\left\{\pm 1\right\}$ Sourlas (1994, 1998, 2001); Sourlas N (2002). Source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$ are mapped to vectors $\boldsymbol{Z}=\left(Z_{1},\ldots,Z_{K}\right)\in\left\{\pm 1\right\}^{K}$ and $\boldsymbol{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ in the spin representation, respectively. Hereafter, we will refer to $\boldsymbol{Z}$ and $\boldsymbol{z}$ as a source-state and a code-state associated with source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$, respectively, and associated fictitious spins as logical and physical spins, respectively. Similarly, binary vectors $\bar{\boldsymbol{s}}$ and $\bar{\boldsymbol{x}}$ are mapped to the associated vectors $\boldsymbol{s}$ and $\boldsymbol{x}$ in the spin representation. In the following, we denote variables by symbols with and without an overbar in the binary and spin representations. It should be noted that by isomorphism, every addition of two binary variables corresponds to a unique product of spin variables and vice versa. For example, since $\bar{\boldsymbol{z}}=\boldsymbol{\bar{Z}}\boldsymbol{G}\>\left(\mathrm{mod}\>2\right)$ holds, $\boldsymbol{z}$ and $\boldsymbol{Z}$ are connected by the relation
@@
for $i=1,\ldots,N_{c}$, since $\bar{\boldsymbol{z}}$ satisfies the parity check equation $\bar{\boldsymbol{z}}\boldsymbol{H}^{T}\>\left(\mathrm{mod}\>2\right)=\boldsymbol{0}_{1\times N_{c}}$.
Now, let us consider the graph in Fig.3, which is topologically equivalent to Fig.2(a). This graph is nothing but the SLHZ model Lechner et al. (2015). This implies that the SLHZ model is essentially a PE-based model that utilizes the weight-4 syndrome, as viewed from the perspective of the spin glass model. The plaquette referred to by LHZ Lechner et al. (2015) and the stabilizer by A. Rocchetto, S. C. Benjamin, and Y. Li Rocchetto et al. (2016) corresponds to the weight-4 syndrome from the perspective of the classical ECCs, each of the elements is given by a product of four elements in $\boldsymbol{x}$ if $\boldsymbol{x}$ is complemented by the fictitious spin variables $x_{ii}\,(i=2,\ldots,K-1)$ with fixed value +1+1: $s_{klmn}^{(4)}(\boldsymbol{x})=x_{km}x_{lm}x_{ln}x_{kn}$ Lechner et al. (2015). The vector $\boldsymbol{s}^{(n)}(\boldsymbol{x})$ should be an all-one vector, i.e., the vector whose elements are all one, for the associated state $\boldsymbol{x}$ of the physical spin to be a valid code-state.
@@ III.1 Probabilistic decoding
We follow the argument given by Sourlas, who first pointed out the link between the ECCs and the spin glass model Sourlas (1989). Sourlas showed that, based on Eqs.(4) and (5), two different formulations are possible for word MAP decoding of the same ECC. The first formulation is given in terms of source-state $\boldsymbol{Z}$. In this formulation, we assume the following prior probability for $\boldsymbol{z}$:
@@
is the half log-likelihood ratio (LLR) for the channel observation $y_{i}$ Sourlas (1994). The vector $\boldsymbol{B}=\left(B_{1},\ldots,B_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ contains all the information about the observation $\boldsymbol{y}$. In this paper, we focus on the communication through the additive white Gaussian noise (AWGN) channel. The likelihood is given as
@@
Alternatively, the second formulation is given in terms of code-state $\boldsymbol{z}$. In this formulation, following Eq.(5), we assume the following prior probability for $\boldsymbol{z}$:
@@
Now, let us recall the SLHZ model. In this case, $s_{i}(\boldsymbol{z})$ is given by the weight-4 syndrome in the spin representation, which is written as a product of four elements in $\boldsymbol{z}$, i.e., $s_{klmn}^{(4)}(\boldsymbol{z})=z_{km}z_{lm}z_{ln}z_{kn}$ Lechner et al. (2015). Then, we see that $H^{code}\left(\boldsymbol{z}\right)$ agrees with the Hamiltonian of the SLHZ model. Therefore, the word MAP decoding of the LDPC codes is equivalent to finding the ground state of the SLHZ model. Furthermore, since the two formulations, i.e., those based on $H^{code}\left(\boldsymbol{z}\right)$ (Eq.(13)) and those based on $H^{source}\left(\boldsymbol{Z}\right)$ (Eq.(8)) are mathematically equivalent, it follows that finding the ground code-state of the SLHZ model is equivalent to finding the ground state of the spin glass.
@@
it represents a metric to measure the reliability or uncertainty of a decision in Eq.(17), which indicates how likely $x_{i}$ it is to be +1+1 or 1-1. A value close to zero indicates a bad inference, while a larger values indicate a better inference. Massey (1962).
A variety of algorithms can achieve symbol MAP decoding. For example, the belief propagation (BP) algorithm is widely regarded as the best-known algorithm for LDPC codes Pearl (1982). The BP algorithm is an iterative, deterministic algorithm in which the a posteriori LLRs, initially given by the observation $\boldsymbol{y}$, gradually increase in absolute value as messages are iteratively exchanged between the VNs and CNs.
In the following section, we study an alternative, simple, and straightforward algorithm: the bit-flipping (BF) algorithm Gallager (1962, 1963). This algorithm is a variety of the symbol MAP decoding strategies. It uses the syndrome vector $\boldsymbol{s}(\boldsymbol{x})=(s_{1}(\boldsymbol{x}),\ldots,s_{N_{c}}(\boldsymbol{x}))$ as a key to decide the most probable value $z_{i}$ of the $i$th spin, namely,
@@ III.2 Readout decoding of the SLHZ model
We first consider the simplest case, i.e., correcting errors caused by i.i.d. noise. This model is appropriate for an SLHZ-based QA device when measurement errors dominate other readout errors. We propose a simple decoding algorithm with very low decoding complexity, which is associated with Gallother algorithm in the spin representation Gallager (1962, 1963). It is based on Massey’s APP (a posteriori probability) algorithm Massey (1962). Our algorithm only uses information about the syndrome vector $\boldsymbol{s}(\boldsymbol{r})$ to estimate errors $\boldsymbol{e}$ in the current decision $\boldsymbol{r}$ for the readout. Let us introduce the error pattern by $\boldsymbol{e}=\boldsymbol{r}\circ\boldsymbol{z}=(e_{12},\ldots,e_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$, where $\circ$ denotes componentwise multiplication. The syndrome vector depends only on $\boldsymbol{e}$ because
@@
Estimating the correct value $z_{ij}$ is equivalent to estimating error $e_{ij}$, as $z_{ij}=r_{ij}e_{ij}$ holds. The algorithm uses the weight-3 syndrome $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=r_{ij}r_{jk}r_{ik}=e_{ij}e_{jk}e_{ik}$ rather than the weight-4 syndrome to decode the current decision $\boldsymbol{r}$ because it has the advantage of treating all variables (VN and CN) symmetrically Pastawski and Preskill (2016). The trade-off is that more constraints may increase the decoding complexity. However, as will be discussed later, this is not a fatal problem, as we can leverage symmetry to reduce computational costs. An optimal estimate $e_{ij}^{*}$ can be obtained from $N_{v}=\tbinom{K}{2}$ syndromes $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)$ for $k\neq i,j$. They are considered K2K-2 parity checks orthogonal on $e_{ij}$ Massey (1962) and their values suggest whether the parity-check equation is satisfied (being +1+1) or violated (being 1-1). Based on the majority voting, we decide an error variable $e_{ij}^{*}$ whether the spin $ij$ should be flipped ($e_{ij}^{*}=-1$) or not ($e_{ij}^{*}=+1$). The optimal estimate $e_{ij}^{*}$ can be obtained according to the APP decoding by Massey (1962)
@@
is called the inversion function Wadayama et al. (2010); Sundararajan et al. (2014), $w_{0}$ and $w_{k}$ are given by
@@
holds Massey (1962); Clerk and Cain (1981). The function $L_{ij}^{*}(\boldsymbol{r})$ can be considered as the LLR for the decision $r_{ij}^{*}$ based on the estimators $\boldsymbol{B}_{ij}(\boldsymbol{r})$, which depends on the weight-3 syndrome vector $\boldsymbol{s}^{(3)}(\boldsymbol{r})$.
Figure 4: Binary symmetric channel defined by the parameters $\gamma_{ij}$ and $p_{k}$.
In Eq.(24), $e_{ij}^{*}=-1$ indicates that we should invert the sign of $r_{ij}\rightarrow r_{ij}^{*}=r_{ij}e_{ij}^{*}=-r_{ij}$. Parameters $\gamma_{ij}$ and $p_{k}$ are crosstalk parameters that characterize the binary symmetric channels shown in Fig.4. We should note that the probability $p_{k}$ is given by the probability of an odd number of 1-1’s among the errors exclusive of $e_{ij}$ that are checked by $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=e_{ij}e_{jk}e_{ik}$ so that it is given by Massey (1962)
@@
Eq.(37) can be interpreted as Gallager’s BF decoding Gallager (1962, 1963). Note that the best estimate $r_{ij}^{*}$ is represented in terms of the values of the current variables in VNs. Fig.5 shows the relevant bipartite graph for $K=5$. Note that the graph associated with the SLHZ system is more loopy, with a minimum length of 4, whereas this graph is less loopy, with a minimum length of 6. The initial decision $\boldsymbol{r}$ on the VNs is broadcast to its adjacent CNs connected by edges. Each CN then reports the syndrome value to its adjacent VNs connected by the edge. All VNs update their values simultaneously according to Eq.(36), representing a majority vote of 1 and the values of adjacent CNs.

### v3

@@ II.2 Parity encoding architecture and LDPC codes
We illustrate a PE architecture with concrete examples. The PE architecture corresponds to the classical ECC according to the following map: $\bar{z}_{ij}=\bar{Z}_{i}\oplus\bar{Z}_{j}$ for $1\leq i\leq j\leq k$. As a simple example, consider the case $K=4$ and assume that $\boldsymbol{\bar{Z}}=\left(\bar{Z}_{1},\bar{Z}_{2},\bar{Z}_{3},\bar{Z}_{4}\right)\in\left\{0,1\right\}^{K}$ and $\bar{\boldsymbol{z}}=\left(\bar{z}_{12},\bar{z}_{13},\bar{z}_{14},\bar{z}_{23},\bar{z}_{24},\bar{z}_{34}\right)\in\left\{0,1\right\}^{\tbinom{K}{2}}$. Because each element of $\bar{\boldsymbol{z}}$ is the binary sum of two elements in $\boldsymbol{\bar{Z}}$, the generating matrix associated with this map is given by the $K\times\tbinom{K}{2}$ matrix:
@@
The connections between the variables $\bar{x}_{kl}$ and the weight-4 syndromes $\bar{s}_{klmn}^{(4)}$ are depicted by a sparse bipartite graph shown in Fig.2(a). Similar connections between variables $\bar{x}_{kl}$ and the weight-3 syndromes $\bar{s}_{klm}^{(3)}$ are depicted in Fig.2(b). The number of elements in $\bar{\boldsymbol{x}}$ is $N_{v}=\tbinom{K}{2}$, while the number of elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ and $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ is $N_{c}=\tbinom{K-1}{2}$ and $N_{c}=\tbinom{K}{3}$, respectively. In the terminology of graph theory, the $N_{v}$ elements in $\bar{\boldsymbol{x}}$ constitute variable nodes (VNs) and the $N_{c}$ elements in $\boldsymbol{\bar{s}}^{(4)}(\bar{\boldsymbol{x}})$ or $\boldsymbol{\bar{s}}^{(3)}(\bar{\boldsymbol{x}})$ constitute check nodes (CNs). The matrix $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$ has $N_{c}$ rows and $N_{v}$ columns, where each row $i$ represents a CN and each column $j$ represents a VN; If the entry $H_{ij}=1$, VN $j$ is connected to CN $i$ by an edge. Thus, each edge connecting VN and CN corresponds to an entry $1$ in the row and column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. Let $d_{c}$ be the number of $1$s in each row and $d_{v}$ be the number of $1$s in each column of $\boldsymbol{H}$ or $\boldsymbol{H}^{\prime}$. These numbers represent the number of edges connected to a VN and a CN, respectively, referred to as row and column weights. The matrix $\boldsymbol{H}^{\prime}$ is regular, where the row weight $d_{c}=3$ is common for all the CNs and the column weight $d_{v}=K-2$ is common for all the VNs. In contrast, the matrix $\boldsymbol{H}$ is irregular, and its weights depend on the associated nodes, which are at most 4.
@@ II.3 SLHZ model
In the remainder of this paper, we will discuss our arguments primarily in the language of spin glass. Following Sourlas, our argument relies on isomorphism between the additive Boolean group $\left(\left\{0,1\right\},\oplus\right)$ and the multiplicative Ising group $\left(\left\{\pm 1\right\},\cdot\right)$, where a binary variable $\bar{a}{}_{i}\in\left\{0,1\right\}$ maps to the spin variable $a_{i}=\left(-1\right)^{\bar{a}_{i}}\in\left\{\pm 1\right\}$ and the binary sum maps to the product by $a_{i}a_{j}=(-1)^{\bar{a}_{i}\oplus\bar{a}_{j}}\in\left\{\pm 1\right\}$ [5, 6, 7, 25]. Source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$ are mapped to vectors $\boldsymbol{Z}=\left(Z_{1},\ldots,Z_{K}\right)\in\left\{\pm 1\right\}^{K}$ and $\boldsymbol{z}=\left(z_{1},\ldots,z_{N_{v}}\right)\in\left\{\pm 1\right\}^{N_{v}}$ in the spin representation, respectively. Hereafter, we will refer to $\boldsymbol{Z}$ and $\boldsymbol{z}$ as a source-state and a code-state associated with source-word $\bar{\boldsymbol{Z}}$ and code-word $\bar{\boldsymbol{z}}$, respectively, and associated fictitious spins as logical and physical spins, respectively. Similarly, binary vectors $\bar{\boldsymbol{s}}$ and $\bar{\boldsymbol{x}}$ are mapped to the associated vectors $\boldsymbol{s}$ and $\boldsymbol{x}$ in the spin representation. In the following, we denote variables by symbols with and without an overbar in the binary and spin representations. It should be noted that by isomorphism, every addition of two binary variables corresponds to a unique product of spin variables and vice versa. For example, since $\bar{\boldsymbol{z}}=\boldsymbol{\bar{Z}}\boldsymbol{G}\>\left(\mathrm{mod}\>2\right)$ holds, $\boldsymbol{z}$ and $\boldsymbol{Z}$ are connected by the relation
@@
for $i=1,\ldots,N_{c}$, if $\bar{\boldsymbol{z}}$ satisfies the parity check equation $\bar{\boldsymbol{z}}\boldsymbol{H}^{T}\>\left(\mathrm{mod}\>2\right)=\boldsymbol{0}_{1\times N_{c}}$.
Now, let us consider the graph in Fig.3, which is topologically equivalent to Fig.2(a). This graph is nothing but the SLHZ model [10]. This implies that the SLHZ model is essentially a PE-based model that utilizes the weight-4 syndrome, as viewed from the perspective of the spin glass model. The plaquette referred to by LHZ [10] and the stabilizer by A. Rocchetto, S. C. Benjamin, and Y. Li [26] corresponds to the weight-4 syndrome from the perspective of the classical ECCs, each of the elements is given by a product of four elements in $\boldsymbol{x}$ if $\boldsymbol{x}$ is complemented by the fictitious spin variables $x_{ii}\,(i=2,\ldots,K-1)$ with fixed value +1+1: $s_{klmn}^{(4)}(\boldsymbol{x})=x_{km}x_{lm}x_{ln}x_{kn}$ [10]. The vector $\boldsymbol{s}^{(n)}(\boldsymbol{x})$ should be an all-one vector, i.e., the vector whose elements are all one, for the associated state $\boldsymbol{x}$ of the physical spin to be a valid code-state.
@@ III.1 Probabilistic decoding
We follow the argument given by Sourlas, who first pointed out the link between the ECCs and the spin glass model [4]. Sourlas showed that, based on Eqs.(4) and (5), two different formulations are possible for word MAP decoding of the same ECC. The first formulation is given in terms of source-state $\boldsymbol{Z}$. In this formulation, we assume the following prior probability for $\boldsymbol{z}$:
@@
is the half log-likelihood ratio (LLR) for the channel observation $y_{i}$ [5]. The vector $\boldsymbol{B}=\left(B_{1},\ldots,B_{N_{v}}\right)\in\mathbb{R}^{N_{v}}$ contains all the information about the observation $\boldsymbol{y}$. In this paper, we focus on the communication through the additive white Gaussian noise (AWGN) channel. The likelihood is given as
@@
Alternatively, the second formulation is given in terms of code-state $\boldsymbol{z}$. In this formulation, according to Eq.(5), we assume the following prior probability for $\boldsymbol{z}$:
@@
Now, let us recall the SLHZ model. In this case, $s_{i}(\boldsymbol{z})$ is given by the weight-4 syndrome in the spin representation, which is written as a product of four elements in $\boldsymbol{z}$, i.e., $s_{klmn}^{(4)}(\boldsymbol{z})=z_{km}z_{lm}z_{ln}z_{kn}$ [10]. Then, we see that $H^{code}\left(\boldsymbol{z}\right)$ agrees with the Hamiltonian of the SLHZ model. Therefore, the word MAP decoding of the LDPC codes is equivalent to finding the ground state of the SLHZ model. Furthermore, since the two formulations, i.e., those based on $H^{code}\left(\boldsymbol{z}\right)$ (Eq.(13)) and those based on $H^{source}\left(\boldsymbol{Z}\right)$ (Eq.(8)) are mathematically equivalent, it follows that finding the ground code-state of the SLHZ model is equivalent to finding the ground state of the spin glass.
@@
it represents a metric to measure the reliability or uncertainty of a decision in Eq.(17), which indicates how likely $x_{i}$ it is to be +1+1 or 1-1. A value close to zero indicates a bad inference, while a larger values indicate a better inference. [27].
A variety of algorithms can achieve symbol MAP decoding. For example, the belief propagation (BP) algorithm is widely regarded as the best-known algorithm for LDPC codes [21]. The BP algorithm is an iterative, deterministic algorithm in which the a posteriori LLRs, initially given by the observation $\boldsymbol{y}$, gradually increase in absolute value as messages are iteratively exchanged between the VNs and CNs.
In the following section, we study an alternative, simple, and straightforward algorithm: the bit-flipping (BF) algorithm [22, 23]. This algorithm is a variety of the symbol MAP decoding strategies. It uses the syndrome vector $\boldsymbol{s}(\boldsymbol{x})=(s_{1}(\boldsymbol{x}),\ldots,s_{N_{c}}(\boldsymbol{x}))$ as a key to decide the most probable value $z_{i}$ of the $i$th spin, namely,
@@ III.2 Readout decoding of the SLHZ model
We first consider the simplest case, i.e., correcting errors caused by i.i.d. noise. This model is appropriate for an SLHZ-based QA device when measurement errors dominate other readout errors. We propose a simple decoding algorithm with very low decoding complexity, which is associated with Gallager’s BF decoding algorithm in the spin representation [22, 23]. It is based on Massey’s APP (a posteriori probability) algorithm [27]. Our algorithm only uses information about the syndrome vector $\boldsymbol{s}(\boldsymbol{r})$ to estimate errors $\boldsymbol{e}$ in the current decision $\boldsymbol{r}$ for the readout. Let us introduce the error pattern by $\boldsymbol{e}=\boldsymbol{r}\circ\boldsymbol{z}=(e_{12},\ldots,e_{K-1\,K})\in\left\{\pm 1\right\}^{\tbinom{K}{2}}$, where $\circ$ denotes componentwise multiplication. The syndrome vector depends only on $\boldsymbol{e}$ because
@@
Estimating the correct value $z_{ij}$ is equivalent to estimating error $e_{ij}$, as $z_{ij}=r_{ij}e_{ij}$ holds. The algorithm uses the weight-3 syndrome $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=r_{ij}r_{jk}r_{ik}=e_{ij}e_{jk}e_{ik}$ rather than the weight-4 syndrome to decode the current decision $\boldsymbol{r}$ because it has the advantage of treating all variables (VN and CN) symmetrically [20]. The trade-off is that more constraints may increase the decoding complexity. However, as will be discussed later, this is not a fatal problem, as we can leverage symmetry to reduce computational costs. An optimal estimate $e_{ij}^{*}$ can be obtained from $N_{v}=\tbinom{K}{2}$ syndromes $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)$ for $k\neq i,j$. They are considered K2K-2 parity checks orthogonal on $e_{ij}$ [27] and their values suggest whether the parity-check equation is satisfied (being +1+1) or violated (being 1-1). Based on the majority voting, we decide an error variable $e_{ij}^{*}$ whether the spin $ij$ should be flipped ($e_{ij}^{*}=-1$) or not ($e_{ij}^{*}=+1$). The optimal estimate $e_{ij}^{*}$ can be obtained according to the APP decoding by [27]
@@
is called the inversion function [28, 29], $w_{0}$ and $w_{k}$ are given by
@@
holds [27, 30]. The function $L_{ij}^{*}(\boldsymbol{r})$ can be considered as the LLR for the decision $r_{ij}^{*}$ based on the estimators $\boldsymbol{B}_{ij}(\boldsymbol{r})$, which depends on the weight-3 syndrome vector $\boldsymbol{s}^{(3)}(\boldsymbol{r})$.
Figure 4: Binary symmetric channel associated with the parameters $\gamma_{ij}$ and $p_{k}$.
In Eq.(24), $e_{ij}^{*}=-1$ indicates that we should invert the sign of $r_{ij}\rightarrow r_{ij}^{*}=r_{ij}e_{ij}^{*}=-r_{ij}$. Parameters $\gamma_{ij}$ and $p_{k}$ are crosstalk parameters that characterize the binary symmetric channels shown in Fig.4. We should note that the probability $p_{k}$ is given by the probability of an odd number of 1-1’s among the errors exclusive of $e_{ij}$ that are checked by $s_{ijk}^{(3)}\left(\boldsymbol{r}\right)=e_{ij}e_{jk}e_{ik}$ so that it is given by [27]
@@
Eq.(37) can be interpreted as Gallager’s BF decoding [22, 23]. Note that the best estimate $r_{ij}^{*}$ is represented in terms of the values of the current variables in VNs. Fig.5 shows the relevant bipartite graph for $K=5$. Note that the graph associated with the SLHZ system is more loopy, with a minimum length of 4, whereas this graph is less loopy, with a minimum length of 6. The initial decision $\boldsymbol{r}$ on the VNs is broadcast to its adjacent CNs connected by edges. Each CN then reports the syndrome value to its adjacent VNs connected by the edge. All VNs update their values simultaneously according to Eq.(36), representing a majority vote of 1 and the values of adjacent CNs.